# Assessing Adaptive World Models in Machines with Novel Games

### v1

@@ 7 Conclusion
In this paper, we have argued that building human-like adaptability in machines necessitates adaptive world models, which affords sample efficient world model induction in any new domains. We then propose a novel framework for assessing adaptive world models centered on the concept of novel games. We believe this proposed evaluation paradigm holds significant potential to serve as a core component in assessing current AI models and drive research towards systems that exhibit the rapid, flexible, and robust adaptability characteristic of human intelligence, thus contributing meaningfully to the ambitious pursuit of artificial general intelligence.

### v2

@@ 7 Conclusion
In this paper, we have argued that building human-like adaptability in machines necessitates adaptive world models, which affords sample efficient world model induction in any new domains. We then proposed a novel framework for assessing adaptive world models centered on the concept of novel games. We believe this proposed evaluation paradigm holds significant potential to serve as a core component in assessing current AI models and drive research towards systems that exhibit the rapid, flexible, and robust adaptability characteristic of human intelligence, thus contributing meaningfully to the ambitious pursuit of artificial general intelligence.